One number is 4 less than a second number. Twice the second number is 1 less than 3 times the first. Find the two numbers.
Answer by mananth(16949)

let second number be x
the first number = x-4
2x =3(x-4)-1
2x=3x-12-1
-x=-13
x= 13
The numbers are 13 & 9